# Taming Rectified Flow for Inversion and Editing

### v2

@@ 3.3 RF-Edit
Incorporating higher-order terms enables RF-Solver to significantly reduce errors in the ODE-solving process, thereby enhancing both sampling quality and inversion accuracy. Furthermore, we extend the application of RF-Solver to the more complex real-world image and video editing tasks, which present greater challenges than reconstruction. In such scenarios, preserving the content and structure of the original image is crucial. For example, when replacing an object in a source image with another one, regions unrelated to the object in this image are expected to remain unaffected by the editing process. However, directly applying RF-Solver during the inversion and denoising stages may cause the model to be overly influenced by the target prompt, resulting in unintended modifications in other regions of the source image or video. Similar issues are common across various existing editing methods [58, 20, 64].
@@
The proposed RF-Edit framework enables high-quality editing while effectively preserving the structural information of the source image/video. Building on this concept, we design two sub-modules for RF-Edit, specifically tailored for image editing and video editing (Fig. 3). For image editing, RF-Edit employs FLUX [2] as the backbone, which comprises several double blocks and single blocks. Double blocks independently modulate text and image features, while single blocks concatenate these features for unified modulation. In this architecture, RF-Edit shares features within the single blocks, as they capture information from both the source image and the source prompt, enhancing the ability of the model to preserve the structural information of the source image. For video editing, RF-Edit employs OpenSora [3] as the backbone. The DiT blocks in OpenSora include spatial attention, temporal attention, and text cross-attention. Within this architecture, the structural information of the source video is captured in the spatial attention module, where we implement feature sharing.
@@ 4.1 Setup
Baselines. We use the vanilla Rectified Flow sampler (Euler Sampler) as the primary baseline for all tasks. Besides, for sampling and inversion, we compare our method with DPM-Solver++ [44] and the Heun sampler (a 2-order ODE solver). For image editing, we compare our method with P2P [20], DiffEdit [11], SDEdit [45], PnP [64], Pix2pix [50] and RF-Inversion [58]. For video editing tasks, we compare our method with FateZero [53], FLATTEN [9], COVE [68], RAVE [30], Tokenflow [16]. Detailed experimental settings of baselines are provided in the Appendix.
Figure 5: Qualitative results of text-to-image generation. By employing the RF-Solver, the model is able to generate images with higher quality (the first row) than baselines (the second row).
Implementation Details. In the experiment, we adopt the guidance-distilled variant of FLUX [2] for image tasks and OpenSora [3] for video tasks. The derivative computation in RF-Solver requires an additional forward pass, resulting in the network needing to forward twice at each timestep. As a result, when comparing our method with the Rectified Flow baselines, we set the number of timesteps for the vanilla Rectified Flow to be twice that of our method to ensure a fair comparison under the same number of function evaluations (NFE). More information is provided in the Appendix.
Evaluation Metrics For text-to-image sampling, we randomly select 10k images from the MSCOCO validation set [39] and report the FID [21] and CLIP Score [54]. For the inversion and reconstruction task, we report the Mean Square Error (MSE), LPIPS [80], SSIM [72], and PSNR [28]. For image editing tasks, we report the CLIP Score [54] and LPIPS [80]. For video editing tasks, we adopt the metric proposed by [26], including Subject Consistency (SC), Motion Smoothness (MS), Aesthetic Quality (AQ), and Imaging Quality (IQ). Detailed explanations of these metrics are provided in the Appendix.
@@ 4.2 Text-to-Image Sampling
We compare the performance of our method with DPMSolver++ [44], the vanilla RF sampler, and Heun sampler on the text-to-image generation task. Both the quantitative (Tab. 1) and qualitative results (Fig. 5) demonstrate the superior performance of RF-Solver in fundamental T2I generation tasks, producing higher-quality images that align more closely with human cognition.
Figure 6: Qualitative comparison of image editing. With RF-Solver and feature-sharing mechanism in RF-Edit, our method can successfully handle various kinds of image editing cases, outperforming the previous SOTA methods. Zoom in for the best views.
DPMSolver++ RF RF-Heun Ours
FID ($\downarrow$) 24.63 25.33 24.40 24.03
CLIP Score ($\uparrow$) 30.62 31.01 31.03 31.09
@@ 4.3 Inversion and Reconstruction
Quantitative Comparison. The quantitative comparisons (Tab. 2) are conducted to reflect the similarity between the source and reconstruction results. Our method demonstrates superior performance across all four metrics compared with the vanilla RF sampler and Heun sampler.
Mehtod MSE ($\downarrow$) LPIPS ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$)
@@
Ours 0.0139 0.3299 0.8805 18.32
Qualitative Comparison. RF-Solver effectively reduces the error in the solution of RF ODE, thereby increasing the accuracy of the reconstruction. As illustrated in Fig. 4(a), the image reconstruction results using vanilla rectified flow exhibit noticeable drift from the source image, with significant alterations to the appearance of subjects in the image. For video reconstruction, as shown in Fig. 4(b), the baseline reconstruction results suffer from distortion. In contrast, RF-Solver significantly alleviates these issues, achieving more satisfactory results.
Figure 7: Qualitative comparison of video editing. The first video comprises 200 frames with a resolution of $512\times 512$, while the second video contains 60 frames with a resolution of $1024\times 768$ (frames are compressed for a neat layout in the figure).
@@ 4.4 Editing
We conduct experiments to evaluate the image and video editing performance of our method. Image editing usually involves replacing the subject in the image with another one, adding new items, and global editing. For the first two types of editing, the background of the source image is expected to remain unchanged after editing. For global editing such as style transfer, the overall structure of the source image is expected to remain unchanged. Recent mainstream video editing methods usually focus on replacing the subjects and performing global editing for the source video.
Quantitative Comparison. In image editing, Our method outperforms all other methods in CLIP score (Tab. 3), indicating that the edited images align well with the user-provided prompts. For LPIPS, it is noted that PnP [64] has a much lower value than all other methods. Based on the qualitative results (Fig. 6), it can be seen that PnP is only suitable for editing cases that do not significantly modify the structure or shape of the source image (such as changing red roses into yellow sunflowers). It fails in the case of shape editing, resulting in an image very similar to the source. Consequently, although PnP has the lowest LPIPS score, its CLIP score is the lowest.
@@
Table 3: Quantitative results of image editing. RF-Edit effectively edit the images according to the prompts while preserving the integrity of unrelated regions.
For video editing, RF-Edit achieves higher scores on the popular VBench [26] metrics (Tab. 4). The results illustrate that our method successfully maintains temporal consistency while demonstrating superior visual quality.
@@
Qualitative Comparison. For image editing, we compare the performance of our method with several baselines across different types of editing tasks (Fig. 6). The baseline methods often suffer from background changes or fail to perform the desired edits. In contrast, our method demonstrates satisfying performance, effectively achieving a balanced trade-off between the fidelity to the target prompt and the preservation of the source image.
For video editing, we primarily evaluate the performance of our method on long videos (200 frames) and high-resolution videos ($1280\times 768$). The qualitative results are shown in Fig. 7. RF-Edit illustrates impressive performance in handling complicated editing cases (e.g., modifying the leftmost lion among three lions into a white polar bear and changing the other two small lions into orange tiger cubs), whereas all other baseline methods fail in this scenario. RF-Edit also demonstrates strong performance in global editing tasks, such as transforming scenes into autumn.
@@
Inversion MSE ($\downarrow$) 0.0268 0.0094 0.0131
LPIPS ($\downarrow$) 0.6253 0.4242 0.4817
@@
Table 5: Ablation study on the Taylor Expansion order. We select the 2-order expansion (i.e. RF-Solver-2) for various downstream tasks due to its effectiveness and simplicity.
4.5 Ablation Study
We conduct ablation studies to illustrate the effectiveness of RF-Solver and RF-Edit. Without loss of generality, these ablation studies are performed on the image tasks using FLUX [2] as the base model.
Taylor Expansion Order of RF-Solver. We investigated the impact of the Taylor expansion order in RF-Solver (Tab. 5) under the same NFE across different orders. The second-order expansion demonstrated a significant improvement across various tasks compared to the first-order expansion (i.e., the vanilla rectified flow). However, higher-order expansions do not yield further enhancements. We speculate that this is primarily due to higher-order Taylor expansions requiring more inference steps per timestep. With a fixed NFE, this results in a reduced overall number of timesteps compared to lower-order expansions, leading to suboptimal performance. Moreover, computing the higher-order derivatives of $\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{t_{i}},{t_{i}})$ substantially increases the complexity of the algorithm, posing challenges for practical applications. Consequently, we predominantly employed second-order expansion in our experiments.
Figure 8: Ablation study of feature-sharing step in RF-Edit. A too-small feature-sharing step results in the inconsistency between source and target images. Conversely, a too-large feature-sharing step can lead to the failure of editing.
Feature Sharing Steps of RF-Edit. RF-Edit leverages feature sharing to maintain the structural consistency between original images and edited images. However, an excessive number of feature-sharing steps may result in the edited output being overly similar to the source image, ultimately undermining the intended editing objectives (Fig. 8). To investigate the impact of feature-sharing steps on editing results, we incrementally increase the number of feature-sharing steps applied to the same image. Due to the varying levels of difficulty that different images presented to the model, the optimal number of sharing steps may differ across cases. Experimental results reveal that setting the sharing step to 5 effectively meets the editing requirements for most images. Additionally, we can customize the sharing step for each image to identify the most satisfying outcome.
@@ 5 Conclusion
Supplementary Material
6 Experimental Settings
6.1 Baselines and Implementation Details
@@
Inversion. We compare the performance of our methods among RF with the vanilla sampler and the Heun sampler. For image inversion, we also use the images from the MS-COCO validation set. For video inversion, we select videos from social media platforms such as TikTok and other publicly available sources. We have observed the quality of the text prompts significantly influence the quality of inversion. Consequently, we employ GPT-4o to generate detailed captions for both images and videos, which are then used in the inversion tasks. The total NFE for generating one image/video is set to 50 for both our method and baselines.
Editing. For image editing, we compare our methods with RF-inversion and several diffusion-based editing methods. For RF-inversion, we adopt the implementation in ComfyUI [1]. For other baselines, we use their implementation from diffusers. For each baseline, we adjust the relevant hyper-parameters to achieve optimal results. For video editing, we use the official codes of all the baseline methods and tune the hyper-parameters to achieve satisfactory results. For image editing, we share the features of the last 19 single blocks in FLUX. For video editing, we share the features of the last 14 blocks in Open-Sora. We adjust the hyper-parameter of feature-sharing steps to achieve better results for both image and video editing.
Figure 9: Stylization Results.
6.2 Evaluation Metrics
@@
For video editing, we adopt the VBench Metrics [26]. The evaluation criteria include Subject Consistency, Motion Smoothness, Aesthetic Quality, and Imaging Quality. Subject Consistency measures whether the subject (e.g., a person) remains consistent throughout the video by computing the similarity of DINO features [6] across frames. Motion Smoothness assesses the smoothness of motion in the generated video using motion priors from the video frame interpolation model [38]. Aesthetic Quality evaluates the artistic and visual appeal of each frame as perceived by humans, leveraging the LAION aesthetic predictor [34]. Imaging Quality examines the level of distortion in the generated frames (e.g., blurring or flickering) based on the MUSIQ image quality predictor [32].
7 Stylization
We provide more results on image stylization (Fig. 9). In this task, we perform inversion on the image with a reference style, following the same overall pipeline as in image editing. RF-Edit demonstrates superior performance in stylized generation, effectively preserving the style of the source image while generating images that align with the user-provided prompts.
8 Limitation
Despite achieving impressive performance across various tasks, our methods occasionally exhibit performance instability, particularly in video editing. Additionally, the FLUX and OpenSora models contain a large number of parameters, leading to significant memory consumption, which may limit the applicability of our method on resource-constrained devices. To mitigate these issues, we plan to further optimize our method and explore its efficacy on more lightweight models.

### v3

@@ 3.3 RF-Edit
Incorporating higher-order terms enables RF-Solver to significantly reduce errors in the ODE-solving process, thereby enhancing both sampling quality and inversion accuracy. Furthermore, we extend the application of RF-Solver to the more complex real-world image and video editing tasks, which present greater challenges than reconstruction. In such scenarios, preserving the content and structure of the original image is crucial. For example, when replacing an object in a source image with another one, regions unrelated to the object in this image are expected to remain unaffected by the editing process. However, directly applying RF-Solver during the inversion and denoising stages may cause the model to be overly influenced by the target prompt, resulting in unintended modifications in other regions of the source image or video. Similar issues are common across various existing editing methods (Rout et al., 2024b; Hertz et al., 2022; Tumanyan et al., 2023).
@@
The proposed RF-Edit framework enables high-quality editing while effectively preserving the structural information of the source image/video. Building on this concept, we design two sub-modules for RF-Edit, specifically tailored for image editing and video editing (Figure 3). For image editing, RF-Edit employs FLUX (Black-Forest-Labs, 2024) as the backbone, which comprises several double blocks and single blocks. Double blocks independently modulate text and image features, while single blocks concatenate these features for unified modulation. In this architecture, RF-Edit shares features within the single blocks, as they capture information from both the source image and the source prompt, enhancing the ability of the model to preserve the structural information of the source image. For video editing, RF-Edit mainly employs OpenSora (Zheng et al., 2024) as the backbone. The DiT blocks in OpenSora include spatial attention, temporal attention, and text cross-attention. Within this architecture, the structural information of the source video is captured in the spatial attention module, where we implement feature sharing.
@@ 4.1 Setup
We implement our method respectively on FLUX (Black-Forest-Labs, 2024) and OpenSora (Zheng et al., 2024). In the experiment, we adopt the guidance-distilled variant of FLUX (Black-Forest-Labs, 2024) for image tasks and OpenSora (Zheng et al., 2024) for video tasks. The derivative computation in RF-Solver requires an additional forward pass, resulting in the network needing to forward twice at each timestep. As a result, when comparing our method with the Rectified Flow baselines, we set the number of timesteps for the vanilla Rectified Flow to be twice that of our method to ensure a fair comparison under the same number of function evaluations (NFE). More detailed information for experiment setup is provided in Appendix B.
DPMSolver++ RF RF-Heun Ours
FID ($\downarrow$) 24.63 25.33 24.40 23.98
CLIP Score ($\uparrow$) 30.62 31.01 31.03 31.09
@@ 4.2 Text-to-image Sampling
We compare the performance of our method with DPMSolver++ (Lu et al., 2022), the vanilla RF sampler, and Heun sampler on the text-to-image generation task. Both the quantitative (Table 1) and qualitative results (Figure 10) demonstrate the superior performance of RF-Solver in fundamental T2I generation tasks, producing higher-quality images that align more closely with human cognition.
Figure 5: Qualitative comparison of image editing. With RF-Solver and feature-sharing mechanism in RF-Edit, our method can successfully handle various kinds of image editing cases, outperforming the previous SOTA methods.
@@ 4.3 Inversion and Reconstruction
Quantitative Comparison. The quantitative comparisons (Table 2) are conducted to reflect the similarity between the source and reconstruction results. Our method demonstrates superior performance across all four metrics compared with the vanilla RF sampler and Heun sampler.
Method MSE ($\downarrow$) LPIPS ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$)
@@
Ours 0.0134 0.3287 0.8812 18.41
Qualitative Comparison. RF-Solver effectively reduces the error in the solution of RF ODE, thereby increasing the accuracy of the reconstruction. As illustrated in Figure 4(a), the image reconstruction results using vanilla rectified flow exhibit noticeable drift from the source image, with significant alterations to the appearance of subjects in the image. For video reconstruction, as shown in Figure 4(b), the baseline reconstruction results suffer from distortion. In contrast, RF-Solver significantly alleviates these issues, achieving more satisfactory results.
@@ 4.4 Editing
We conduct experiments to evaluate the image and video editing performance of our method. For image editing, we compare our method with P2P (Hertz et al., 2022), DiffEdit (Couairon et al., 2023), SDEdit (Meng et al., 2022), PnP (Tumanyan et al., 2023), Pix2pix (Parmar et al., 2023) and RF-Inversion (Rout et al., 2024b). For video editing tasks, we compare our method with FateZero (Qi et al., 2023), FLATTEN (Cong et al., 2023), COVE (Wang et al., 2024c), RAVE (Kara et al., 2024), Tokenflow (Geyer et al., 2023).
Table 3: Quantitative results of image editing. RF-Edit effectively edits the images according to the prompts while preserving the integrity of unrelated regions.
@@
Quantitative Comparison. In image editing, Our method outperforms all other methods in CLIP score (Table 3), indicating that the edited images align well with the user-provided prompts. For LPIPS, it is noted that PnP (Tumanyan et al., 2023) has a much lower value than all other methods. Based on the qualitative results (Figure 5), it can be seen that PnP is only suitable for editing cases that do not significantly modify the structure or shape of the source image (such as changing red roses into yellow sunflowers). It fails in the case of shape editing, resulting in an image very similar to the source. Consequently, although PnP has the lowest LPIPS score, its CLIP score is the lowest. For video editing, RF-Edit achieves higher scores on VBench (Huang et al., 2024b) metrics (Table 4). The results illustrate that our method successfully maintains temporal consistency while demonstrating superior quality.
Figure 6: Qualitative comparison of video editing. The first video comprises 200 frames with a resolution of $512\times 512$, while the second video contains 60 frames with a resolution of $1024\times 768$ (frames are compressed for a neat layout).
@@
Qualitative Comparison. For image editing, we compare the performance of our method with several baselines across different types of editing requirements including adding, replacing, and stylization (Figure 5 and Figure 8). The baseline methods often suffer from background changes or fail to perform the desired edits. In contrast, our method demonstrates satisfying performance, effectively achieving a balanced trade-off between the fidelity to the target prompt and the preservation of the source image.
The qualitative results for video editing are shown in Figure 6. RF-Edit illustrates impressive performance in handling complicated editing cases (e.g., modifying the leftmost lion among three lions into a white polar bear and changing the other two small lions into orange tiger cubs), whereas all other baseline methods fail in this scenario.
Besides, HunyuanVideo (Kong et al., 2024) has recently demonstrated strong performance in text-to-video generation. Thanks to the generality of our method, it can be implemented on HunyuanVideo for video editing. More qualitative results are shown in Figure 9.
4.5 Ablation Study
We conduct ablation studies to illustrate the effectiveness of RF-Solver and RF-Edit. Without loss of generality, these ablation studies are performed on the image tasks using FLUX (Black-Forest-Labs, 2024) as the base model.
Taylor Expansion Order of RF-Solver. We investigated the impact of the Taylor expansion order in RF-Solver (Table 5) under the same NFE across different orders. Compared to the first-order expansion (i.e., the vanilla rectified flow), the second-order expansion demonstrates a significant improvement across various tasks. However, higher-order expansions do not yield further enhancements. We speculate that this is primarily due to higher-order Taylor expansions requiring more inference steps per timestep. With a fixed NFE, this results in a reduced overall number of timesteps compared to lower-order expansions, leading to suboptimal performance. Moreover, computing the higher-order derivatives of $\boldsymbol{v}_{\theta}(\boldsymbol{Z}_{t_{i}},{t_{i}})$ substantially increases the complexity of the algorithm, posing challenges for practical applications. Consequently, we employ second-order expansion (i.e., RF-Solver-2 in Table 5) for various downstream tasks due to its effectiveness and simplicity.
Table 5: Ablation study on the Taylor Expansion order. Here, RF implies the vanilla Rectified Flow without the proposed RF-Solver algorithm. All of the experiments for editing in the table apply the proposed feature-sharing mechanism for better results.
@@ 4.5 Ablation Study
Inversion MSE ($\downarrow$) 0.0268 0.0092 0.0131
LPIPS ($\downarrow$) 0.6253 0.4239 0.4817
@@
Feature Sharing Steps of RF-Edit. RF-Edit leverages feature sharing to maintain the structural consistency between original images and edited images. However, an excessive number of feature-sharing steps may result in the edited output being overly similar to the source image, ultimately undermining the intended editing objectives (Figure 7). To investigate the impact of feature-sharing steps on editing results, we incrementally increase the number of feature-sharing steps applied to the same image. Due to the varying levels of difficulty that different images presented to the model, the optimal number of sharing steps may differ across cases. Experimental results reveal that setting the sharing step to 5 effectively meets the editing requirements for most images. Additionally, we can customize the sharing step for each image to identify the most satisfying outcome.
Figure 7: Ablation study of feature-sharing step in RF-Edit. The second row is the results produced solely by RF-Solver, without the proposed feature-sharing mechanism. Feature-sharing can significantly enhance the consistency between source and target images, while a too-large feature-sharing step may lead to the failure of editing.
@@ 5 Conclusion
Acknowledgments
This work was partly supported by Shenzhen Key Labora-tory of next generation interactive media innovative tech-nology (No:ZDSYS20210623092001004) and National Key Laboratory of Human-Machine Hybrid Augmented Intelligence, Xi’an Jiaotong University (No. HMHAI202410).
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix
Appendix A Pesudo Code of RF-Solver Algorithm.
Algorithm 1 Sampling process of RF-Solver
Input:
  $\boldsymbol{v}_{\theta}$triangleright Velocity function
  $t=[t_{N},\ldots,t_{0}]$triangleright Time steps
For $i=N$ to $1$ do
Output: $\boldsymbol{Z}_{0}$
Appendix B Experimental Settings
B.1 Baselines and Implementation Details
@@
Inversion. We compare the performance of our methods among RF with the vanilla sampler and the Heun sampler. For image inversion, similar to Text-to-Image generation, we use images and captions from the MS-COCO validation set. For video inversion, we select about 40 videos from social media platforms such as TikTok and other publicly available sources. We have observed the quality of the text prompts significantly influences the quality of inversion. Consequently, we employ GPT-4o to generate detailed captions for both images and videos, which are then used in the inversion tasks. The total NFE for generating one image/video is set to 50 for both our method and baselines.
Figure 8: Stylized Generation.
Figure 9: More video editing results.
Editing. For image editing, we share the features of the last 19 single blocks in FLUX. For video editing, we share the features of the last 14 blocks in Open-Sora. We adjust the hyper-parameter of feature-sharing steps to achieve better results for both image and video editing. For image editing, we use over 300 images for quantitative comparison, which both include real images (obtained from public social media and the DIV2K dataset) and generated images. For each image, we use GPT-4o to generate the source prompt and manually refine the generated prompt. There are 2 3̃ target prompts for different requirements of editing including adding, replacing, and stylization for each source image. We compare our methods with RF-inversion and several diffusion-based editing methods. For RF-inversion, we adopt the implementation in ComfyUI (com, ). For other baselines, we use their implementation from diffuser and adjust the relevant hyper-parameters to achieve optimal results. For video editing, the data preparation mainly follows previous works COVE (Wang et al., 2024c). We use the official codes of all the baseline methods and tune the hyper-parameters to achieve satisfactory results.
Figure 10: Qualitative results of text-to-image generation. By employing the RF-Solver, the model is able to generate images with higher quality than baselines.
B.2 Evaluation Metrics
@@
For video editing, we adopt the VBench Metrics. The evaluation criteria include Subject Consistency, Motion Smoothness, Aesthetic Quality, and Imaging Quality. Subject Consistency measures whether the subject (e.g., a person) remains consistent throughout the video by computing the similarity of DINO features (Caron et al., 2021) across frames, which is similar to the CLIP Score for images. Motion Smoothness assesses the smoothness of motion in the generated video using motion priors from the video frame interpolation model (Li et al., 2023). Aesthetic Quality evaluates the artistic and visual appeal of each frame as perceived by humans, leveraging the LAION aesthetic predictor (LAION-AI, 2022). Imaging Quality examines the level of distortion in the generated frames (e.g., blurring or flickering) based on the MUSIQ image quality predictor (Ke et al., 2021).
Appendix C Qualitative Results for Text-to-image Sampling
Qualitative results for Text-to-image Sampling are shown in Figure 10. Experimental results demonstrate that compared to vanilla Rectified Flow, the RF-Solver sampler can generate higher-quality images that better align with human perception.
Appendix D More Qualitative Results for Image Editing
Here we provide more qualitative results for image editing and stylization (Figure 8).
Appendix E More Qualitative Results for Video Editing
HunyuanVideo (Kong et al., 2024) has recently demonstrated strong performance in text-to-video generation. The backbone of HunyuanVideo is similar to FLUX, which also contains several double-stream blocks, followed by single-stream blocks. We implement the RF-Solver and RF-Edit on HunyuanVideo, where the RF-Edit shares the feature in the single-stream block. The results are shown in Figure 9.
Appendix F More Potential Applications
RF-Solver is a universal sampler for rectified flow. Besides image and video editing, it is also potential on image or video generation (Xiao et al., 2025; Ma et al., 2024b, a; Lin et al., 2025) and other diffusion-based tasks (He et al., 2024; Fang et al., 2024; Guo et al., 2022). Furthermore, our proposed feature-sharing method in RF-Edit can also be applied in other image and video editing methods (Zhu et al., 2024b, a).